Asymptote...I need to find the vertical asymptote of the rational function
f(x)=3x-12/4x-2. I have tried and searched and searched!

Answer by stanbon(75887)

Asymptote...I need to find the vertical asymptote of the rational function
f(x)=3x-12/4x-2
---------
You have a vertical asymptote when the denominator is zero.
So, find out when the denominator is zero, as follows:
4x-2=0
4x=2
x=1/2
You have a vertical asymptote at x=1/2
Cheers,
Stan H.